OpenStudy (anonymous):

Find the roots of this equation 9x^2+12x+4=0

OpenStudy (nikita2):

D = 36 - 9*4 = 0 So x = -6/9 = -2/3

OpenStudy (anonymous):

how do you do it by hand calcs?

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

that results in -12 +or - sqrt 0 the quadratic cant be used can it

OpenStudy (anonymous):

what's wrong with that?

OpenStudy (anonymous):

using the quadratic formula, you end up with -12/18=-2/3

OpenStudy (anonymous):

yeah sorry then divide by 2a gives -2/3 thanks

OpenStudy (anonymous):

no problem